# Attributed-graphs kernel implementation using local detuning of neutral-atoms Rydberg Hamiltonian

### v1

@@ III.2 Generalized-distance quantum-correlation (GDQC) kernel
The kernel function $\kappa_{\mathrm{GDQC}}$ defined in Eq. (13) is: (1) A kernel function (symmetric positive semi-definite), (2) valid on graph data for any graph distance $d_{\mathcal{G}}$, and (3) equal to 1 on any isomorphic pair $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$.
Next, we demonstrate that the GDQC kernel function $\kappa_{\mathrm{GDQC}}$ (13) satisfies all graph-kernel properties. Its key properties are summarized in proposition 2, and are proven in detail in App. B. Let us emphasize that the converse of point (3) is not proven to be true, that is, we do not claim that $\kappa_{\mathrm{GDQC}}(\mathcal{G}_{1},\mathcal{G}_{2})=1$ if and only if $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ are isomorphic. A proof thereof (if it was true) would require a deeper analysis.
The similarity between the introduced GDQC kernel (13) and the Weisfeiler-Leman kernel [6], more precisely its generalized-distance Weisfeiler-and-Leman (GD-WL) version, can be made more explicit. To establish such a comparison, we first need to recall the notion of expressiveness in the context of kernel functions. Let $\kappa_{1}$ and $\kappa_{2}$ be two normalized graph kernels.
@@
    A graph kernel $\kappa$ is expressive for a pair of two non-isomorphic graphs $\mathcal{G}$ and $\mathcal{G}^{\prime}$ if $\kappa(\mathcal{G},\mathcal{G}^{\prime})<1$.
@@
    A kernel $\kappa_{1}$ is more expressive than a kernel $\kappa_{2}$ if for any pair of non-isomorphic graphs, $\mathcal{G}$ and $\mathcal{G}^{\prime}$, two conditions are satisfied. First, $\kappa_{1}(\mathcal{G},\mathcal{G}^{\prime})=1$ always implies $\kappa_{2}(\mathcal{G},\mathcal{G}^{\prime})=1$. Second, there exists at least one pair of non-isomorphic graphs, $\mathcal{G}_{*}$ and $\mathcal{G}^{\prime}_{*}$, such that $\kappa_{2}(\mathcal{G}_{*},\mathcal{G}^{\prime}_{*})=1$ but $\kappa_{1}(\mathcal{G}_{*},\mathcal{G}^{\prime}_{*})<1$.
Here, a practical question concerns the choice of parameters that determine the lengths of the intervals $\Delta_{i}^{(d)}$ and $\Delta_{i}^{(q)}$. Should one favor a larger or smaller number of bins ? Intuitively, each approach has its advantages and drawbacks. The advantage comes mainly from the fact that increasing the number of bins increases the expressiveness of our kernels, which can be characterized by the following property:
@@
The kernel function $\kappa_{\mathrm{GDQC}}$ is as expressive as the GD-WL test [37] given the latter is initialized with colors obtained from the multi-set of the values of the correlation matrix rows/columns.
A proof of this proposition is provided in appendix C. One can also point out that if the selected graph distance is such that the corresponding GD-WL is less expressive than the quantum correlation matrix, then $\kappa_{\mathrm{GDQC}}$ expressiveness is upper bounded by the expressiveness of the quantum correlation matrix ($QC$), as it already provides a stable coloring for the GD-WL (more details about GD-WL and its node coloring are to be found in Appendix C). This point is crucial, as the authors in [38] show that for several graph distances such as Laplacian eigenvectors, resistance distance and shortest path distance, GD-WL is upper bounded by 3-WL. On the other hand, the authors in [10] empirically show that the quantum correlation matrix of local observables issued from the Ising Hamiltonian is at least as powerful as 4-WL, as they are able to distinguish strongly regular graphs [39, 40]. When combined, these results provide empirical evidence that $\kappa_{\mathrm{GDQC}}$ is upper bounded by $QC$ for the state-of-the-art graph distances studied in [38]. This suggests that computing $\kappa_{\mathrm{GDQC}}$ with "simple" graph distances (ex. shortest path distance) does not improve the kernel’s expressiveness, but it does not necessarily suggest that it does not improve the performances of the subsequent machine learning model (cf. Sec. IV). Indeed, the main inconvenient for raising the number of bins (and thus improving the expressiveness) is that the corresponding Gram matrix would converge towards a binary matrix, which is not a suitable feature for kernel-based machine learning models. In fact, if one has a data set $\mathcal{D}$ for which the kernel is fully expressive, and if this dataset contains no isomorphic pair, then the Gram matrix would simply be the identity matrix, and one knows that such an input does not provide useful information for any machine learning model.
III.3 Kernel pooling
In this part we describe a scheme that allows to combine information that comes from observables sampled at multiple time steps $t_{1},\dots,t_{N_{t}}$}, cf. Eq. (1). In principle, two qualitatively different strategies, i.e. two conceptually different functions $F_{\kappa}$, are conceivable. Either one feeds a large quantum-feature vector $\bm{O}^{\mathcal{G}}$ comprising all time steps into a single kernel function, or one feeds all single-time-step quantum-feature vectors into separate kernel functions which are subsequently pooled. We proceed according to the second strategy, i.e. we generate a collection of $N_{t}$ kernel functions $\kappa_{1},\dots,\kappa_{N_{t}}$}. Let $f:\mathbb{R}^{N_{t}}\to\mathbb{R}$ be a function that preserves positive semi-definiteness (PSD), then by Mercer’s theorem
$\kappa(\mathcal{G}_{1},\mathcal{G}_{2})$ $=f\left[\kappa_{1}(\mathcal{G}_{1},\mathcal{G}_{2}),...,\kappa_{N_{t}}(\mathcal{G}_{1},\mathcal{G}_{2})\right],$ (14)
the resulting expression $\kappa(\mathcal{G}_{1},\mathcal{G}_{2})$ is also a valid kernel function. This includes polynomial kernels, exponential kernels and RBF-like kernels. In our case, we employ two pooling schemes: positive (convex) linear combinations defined by
$\kappa_{+}(\mathcal{G}_{1},\mathcal{G}_{2})=\sum_{\gamma=1}^{N_{t}}\alpha_{\gamma}\kappa_{\gamma}(\mathcal{G}_{1},\mathcal{G}_{2})$ (15)
where $\sum_{\gamma}\alpha_{\gamma}=1$, and an element-wise product pooling defined by
$\kappa_{\times}(\mathcal{G}_{1},\mathcal{G}_{2})=\prod_{\gamma=1}^{N_{t}}\kappa_{\gamma}(\mathcal{G}_{1},\mathcal{G}_{2}).$ (16)
From another perspective, this translates into the following : the feature map $\chi(x)$ corresponding to the kernel $\kappa$ can be expressed in terms of the feature maps $\chi_{i}$ of the kernels $\kappa_{i}$. In the case of Eq. (15) we have :
$\chi_{+}(\mathcal{G})=||_{\gamma=1}^{N_{t}}(\sqrt{\alpha_{\gamma}}\chi_{\gamma}(\mathcal{G}))$ (17)
where $||_{\gamma=1}^{N_{t}}$ is the concatenation operator, and in the case of Eq. (16) we have
$\chi_{\times}(\mathcal{G})=\bigotimes_{\gamma=1}^{N_{t}}\chi_{\gamma}(\mathcal{G})$ (18)
where $\bigotimes$ is the tensor product operator. This allows us to state the intuitive result formulated in proposition 4, see App. D for a detailed proof. It should be emphasized that according to proposition 4 it is possible that pooling improves the expressiveness, yet, it is not guaranteed. It depends on the expressiveness of the added kernels.
@@ III.3 Kernel pooling
For a set $S=\{\kappa_{1},\kappa_{2},...,\kappa_{N}\}$ of graph kernels, let $\kappa_{A}$ and $\kappa_{B}$ be respectively the kernels obtained from pooling two subsets $S_{A}$ and $S_{B}$ of $S$. If $S_{A}\subset S_{B}$ then $\kappa_{A}$ is at least as expressive as $\kappa_{B}$.
@@
Classical
Random Walk 84.37 $\pm$ 5.80 52.33 $\pm$ 4.17
Subgraph sampling 81.70 $\pm$ 6.18 57.25 $\pm$ 4.55
$k$-core 82.84 $\pm$ 5.24 59.10 $\pm$ 5.16
Shortest path 80.72 $\pm$ 5.99 59.92 $\pm$ 4.97
WL optimal assignment 84.64 $\pm$ 5.12 58.34 $\pm$ 5.10
$\text{$k$-core}^{\mathrm{(attr)}}$ 81.69 $\pm$ 6.13 59.10 $\pm$ 5.16
$\text{Shortest path}^{\mathrm{(attr)}}$ 80.00 $\pm$ 6.57 59.95 $\pm$ 4.64
$\text{WL optimal assignment}^{\mathrm{(attr)}}$ 86.10 $\pm$ 5.11 64.96 $\pm$ 5.18
Quantum features
$\mathrm{QEK}^{\mathrm{(loc)}}$ 84.64 $\pm$ 5.05 60.87 $\pm$ 5.34
$\mathrm{QEK}^{\mathrm{(glob)}}$ 84.57 $\pm$ 5.25 61.06 $\pm$ 5.37
$\mathrm{GDQC}^{\mathrm{(loc)}}$ 86.78 $\pm$ 4.53 58.79 $\pm$ 5.39
$\mathrm{GDQC}^{\mathrm{(glob)}}$ 86.09 $\pm$ 4.97 56.92 $\pm$ 4.79
$\mathrm{QEK}^{\mathrm{(loc)}}_{\bm{+}}$ 87.08 $\pm$ 5.81 62.40 $\pm$ 5.77
$\mathrm{QEK}^{\mathrm{(loc)}}_{\bm{\times}}$ 87.41 $\pm$ 6.20 63.89 $\pm$ 6.15
$\mathrm{QEK}^{\mathrm{(glob)}}_{\bm{+}}$ 86.33 $\pm$ 5.10 62.74 $\pm$ 5.12
$\mathrm{QEK}^{\mathrm{(glob)}}_{\bm{\times}}$ 86.53 $\pm$ 4.80 62.42 $\pm$ 5.66
$\mathrm{GDQC}^{\mathrm{(loc)}}_{\bm{+}}$ 87.30 $\pm$ 5.10 62.18 $\pm$ 5.20
$\mathrm{GDQC}^{\mathrm{(loc)}}_{\bm{\times}}$ 86.88 $\pm$ 4.90 65.38 $\pm$ 5.40
$\mathrm{GDQC}^{\mathrm{(glob)}}_{\bm{+}}$ 88.06 $\pm$ 5.00 59.22 $\pm$ 5.80
$\mathrm{GDQC}^{\mathrm{(glob)}}_{\bm{\times}}$ 87.77 $\pm$ 4.72 60.33 $\pm$ 5.50
@@ IV Results
In this section we present the extensive numerical results where we compare the various classical and quantum-feature kernel implementations introduced in Sec. III with respect to their prediction performance using a support vector machine (SVM). Table 1 summarizes the results, providing a full comparison across all models and their respective best performances.
@@
Throughout this work the precomputed kernels are used to train a SVM on a classification task, see beginning of Sec. III for more details on kernel-based SVMs. The classification performance is evaluated using a weighted F1 score, a classification metric designed to address class imbalances in datasets such as in MUTAG and PTC_FM. The weighted F1 score calculates the F1 score (the harmonic mean of precision and recall) for each class and then takes the weighted average based on the number of instances in each class. Here, we employ a repeated stratified 5-fold cross-validation with 10 repetitions, and we report the mean and standard deviation of the weighted F1 score across all cross-validation splits and repetitions. Furthermore, for every single kernel we perform a hyper-parameter tuning of the regularization parameter $C$ of a SVM. We tune over 100 logarithmically spaced values in the interval $C\in[0.001,100]$, and proceed with the best-performing SVM classifier. The computations involving the classical kernels conducted using the GraKel library [44]. The respectively best results are presented in Tab. 1. Here, one notes that taking into account the node features generally improves the results. Moreover, the best performance is achieved by the WL optimal assignment kernel with supplied node features.
@@
Figure 3: Weighted F1 scores associated with SVMs trained on the quantum-feature kernels, QEK (12) and GDQC (13), for the two datasets, (a) MUTAG and (b) PTC_FM. The superscripts $\mathrm{(loc)}$ and $\mathrm{(glob)}$ refer to whether the governing Hamiltonian has node features embedded $\hat{\mathcal{H}}_{\mathrm{loc}}^{\mathcal{G}}$ (10)] or not $\hat{\mathcal{H}}_{\mathrm{glob}}^{\mathcal{G}}$ (9)]. The training and evaluation procedure is separately conducted at each time steps across the time evolution. For comparison, the corresponding randomly stratified baseline, where each class is predicted proportionally to its frequency in the dataset are 0.55 for MUTAG and 0.51 for PTC_FM, significantly below the presented F1 scores.
Next, we focus on the quantum-feature kernels. The observables are sampled at $k=20$ equally-spaced time steps $\{t_{1},t_{2},...,t_{k}=t_{\mathrm{max}}\}$ and respectively, two emulations (or time evolutions) are conducted, a shorter one with $t_{\mathrm{max}}=1\,\mathrm{\mu s}$ and a longer one with $t_{\mathrm{max}}=2\,\mathrm{\mu s}$. We begin with non-pooled results. At each time step a quantum-feature kernel, $\mathrm{QEK}^{\mathrm{(loc/glob)}}$ or $\mathrm{GDQC}^{\mathrm{(loc/glob)}}$, is precomputed and used to train a SVM. The resulting F1 scores for $t_{\mathrm{max}}=1\,\mathrm{\mu s}$ and for both datasets, MUTAG and PTC_FM, are presented in Fig. 3. Overall, both kernels, QEK and GDQC, yield comparative results; importantly, all results are significantly above the random baseline (black dashed line). In terms of the differences in using the node-feature embedded Hamiltonian $\hat{\mathcal{H}}_{\mathrm{loc}}^{\mathcal{G}}$ [superscript $\mathrm{(loc)}$] or the global-detuning Hamiltonian $\hat{\mathcal{H}}_{\mathrm{glob}}^{\mathcal{G}}$ [superscript $\mathrm{(glob)}$], no clear trend can be observed. As a function of the sampling time steps, the performances vary slightly within a range of $\sim 5\%$ in F1 score. The maxima in the sequence, i.e. the best-performing models are added to Tab. 1. Interestingly, for MUTAG the GDQC performance is on par with the best classical score whereas for PTC_FM the QEK kernel has achieved a higher score than all classical models except for the WL optimal $\mathrm{assignment^{(attr)}}$.
Next, we employ pooling schemes in order to potentially further enhance the performance of the quantum-feature kernels. The two pooling scheme which we apply are described in Subsec. III.3. However, the decision on which of the $k=20$ time steps to pool, and how many time steps to use is quite arbitrary. Instead of relying on a predetermined strategy, we fix the number of time steps to use, and randomly select $1000$ tuples of time-steps combinations. The goal of this study is really to comprehend by how much the performance of our models can be enhanced by means of information pooling. For this reason, we only report the best-performing models for each tuple size.
Having chosen the tuple sizes to $3$ and $5$, we show the best-performing models for $t_{\mathrm{max}}=1\,\mathrm{\mu s}$ and $t_{\mathrm{max}}=2\,\mathrm{\mu s}$ in Tab. 2. For each model we list the involved pooled time steps. The first thing to notice is that pooling can significantly enhance the model performance. With the best-performing pooling models from each kernel class added to Tab. 1 we achieve an F1-score increase by $\sim 2\%$ on the MUTAG dataset and by $\sim 4\%$ for PTC_FM. As a result, we find the best-performing quantum-feature models outperforming the classical algorithms. Considering the results in Tabs. 1-2 one notices a fairly common trend which is a high F1 score on one of the datasets, MUTAG or PTC_FM, being accompanied by a low score on the other dataset. This is particularly pronounced for the classical random walk and shortest-distance kernels. Similarly, $\mathrm{GDQC}^{\mathrm{(glob)}}_{\bm{+}}$ achieves the highest F1 scores on the MUTAG dataset across all models, but it scores among the lowest on PTC_FM, see Tab. 2. Albeit not being statistically significant, in Tab. 2 we find $5$ out of the top $6$ best-performing models associated with the GDQC kernel, maybe hinting at a potential edge of GDQC over the QEK kernel.

### v2

@@ III.3 Generalized-distance quantum-correlation ($\mathrm{GDQC}$) kernel
The function $\kappa_{\mathrm{GDQC}}$ defined in Eq. (17) is a valid graph kernel, i.e. it is symmetric positive semi-definite on the space of graphs equipped with any graph distance $d_{\mathcal{G}}$, and is isomorphism-invariant such that $\kappa_{\mathrm{GDQC}}(\mathcal{G}_{\mu},\mathcal{G}_{\mu^{\prime}})=1$ whenever $\mathcal{G}_{\mu}$ and $\mathcal{G}_{\mu^{\prime}}$ are isomorphic.
Next, we demonstrate that the $\mathrm{GDQC}$ kernel function $\kappa_{\mathrm{GDQC}}$ (17) satisfies all graph-kernel properties. Its key properties are summarized in proposition 2, and are proven in detail in App. D. Let us emphasize that the converse of point (3) is not proven to be true, that is, we do not claim that $\kappa_{\mathrm{GDQC}}(\mathcal{G}_{\mu},\mathcal{G}_{\mu^{\prime}})=1$ if and only if $\mathcal{G}_{\mu}$ and $\mathcal{G}_{\mu^{\prime}}$ are isomorphic. A proof thereof (if it was true) would require a deeper analysis.
The connection between the proposed $\mathrm{GDQC}$ kernel (17) and the Weisfeiler-Lehman (WL) family of kernels [47], in particular its generalized-distance variant (GD-WL), can be made explicit by comparing their expressiveness. Let $\kappa_{1}$ and $\kappa_{2}$ be two normalized graph kernels.
@@
    A kernel $\kappa$ is expressive for a pair of non-isomorphic graphs $(\mathcal{G},\mathcal{G}^{\prime})$ if it distinguishes them, i.e., $\kappa(\mathcal{G},\mathcal{G}^{\prime})<1$.
@@
    We say that $\kappa_{1}$ is more expressive than $\kappa_{2}$ if (i) for all non-isomorphic $(\mathcal{G},\mathcal{G}^{\prime})$, $\kappa_{1}(\mathcal{G},\mathcal{G}^{\prime})=1\Rightarrow\kappa_{2}(\mathcal{G},\mathcal{G}^{\prime})=1$, and (ii) there exists at least one non-isomorphic pair $(\mathcal{G}_{*},\mathcal{G}^{\prime}_{*})$ such that $\kappa_{2}(\mathcal{G}_{*},\mathcal{G}^{\prime}_{*})=1$ but $\kappa_{1}(\mathcal{G}_{*},\mathcal{G}^{\prime}_{*})<1$.
A practical question then concerns the binning parameters $N_{\mathrm{bins}}^{D}$ and $N_{\mathrm{bins}}^{C}$, which determine the bin widths $\Delta_{D}$ and $\Delta_{C}$. Increasing the number of bins refines the induced features and typically increases expressiveness, as formalized by Proposition 3 whose proof is provided in App. E.
@@
The $\mathrm{GDQC}$ kernel $\kappa_{\mathrm{GDQC}}$ matches the expressiveness of the GD-WL refinement [54] when GD-WL is initialized with node colors constructed from the multisets of (binned) correlation values in the rows/columns of $C^{\mathcal{G}}(t)$.
If the chosen graph distance induces a GD-WL refinement that is less expressive than the quantum correlation matrix, then the expressiveness of $\kappa_{\mathrm{GDQC}}$ cannot exceed that of the correlation matrix, since the latter already provides a stable initial coloring for GD-WL (see Appendix E). This point is crucial: Zhang et al. [55] show that, for several commonly used distances (including Laplacian-eigenvector, resistance, and shortest-path distances), the expressiveness of GD-WL is upper bounded by $3$-WL. In contrast, Thabet et al. [51] provide evidence that correlation matrices of local observables generated by an Ising Hamiltonian can be at least as powerful as $4$-WL, as demonstrated by their ability to distinguish strongly regular graphs [21, 8]. Taken together, these results suggest that, for the state-of-the-art distances studied in [55], using a “simple” distance (e.g., shortest-path distance) is unlikely to increase the expressiveness of $\kappa_{\mathrm{GDQC}}$ beyond what is already captured by the quantum correlations, although it may still improve downstream predictive performance (cf. Sec. IV).
Finally, we provide intuition for the $\mathrm{GDQC}$ kernel by identifying a purely classical baseline. To this end, we exploit the early-time dynamics discussed in Subsec. II.4. We have shown that the diagonal correlation-matrix elements scale as $C_{i,i}\left(t\right)=\mathcal{O}\left((\Omega t)^{2}\right)$, whereas the off-diagonal elements scale as $C_{i,j\neq i}\left(t\right)=\mathcal{O}\left((\Omega t)^{4}\right)$ for $i\neq j$. Consequently, at sufficiently early times all entries of $C(t)$ are small. The $\mathrm{GDQC}$ correlation binning therefore assigns all $N^{2}$ node pairs to the first correlation bin, while the remaining bins are empty. Hence, the intermediate correlation-matrix binning vector becomes
$\bm{\chi}^{C}_{\mathrm{bin},\,\mathrm{baseline}}$ $=\left(N^{2},0,\dots,0\right)^{\intercal},$ (18)
cf. Fig. 2. In the near-zero correlation regime, the correlation-binning step carries no discriminative information, so the subsequent distance binning becomes the only nontrivial component. Consequently, the $\mathrm{GDQC}$ feature vector $\bm{\chi}$ is dominated by the graph’s distance profile: for each shortest-path length, it counts (up to normalization) how many node pairs lie at that distance, making its early-time behavior closely related to the classical (unattributed) shortest-path kernel (see Subsec. III.1). Notably, the representation is not strictly identical, since $\mathrm{GDQC}$ also includes contributions from node pairs at zero graph distance, whereas the standard shortest-path kernel excludes diagonal terms. Overall, the early-time $\mathrm{GDQC}$ regime provides a natural classical baseline.
III.4 Kernel pooling
We now describe how to combine information from observables sampled at multiple times $\{t_{1},\dots,t_{N_{t}}\}$ (cf. Eq. 1), while still keeping relevant input for the task at hand. Two strategies are possible: (i) concatenate all time-dependent features into a single vector and apply one kernel, or (ii) define one kernel per time step and pool these kernels. We follow the second strategy and construct a collection of kernels $\{\kappa_{1},\dots,\kappa_{N_{t}}\}$, where $\kappa_{\gamma}$ is computed from observables at time $t_{\gamma}$.
Any pooling rule that preserves kernel properties [cf. proposition 2] yields a valid kernel. In particular, if each $\kappa_{\gamma}$ is a valid kernel, then (a) any convex combination and (b) the pointwise product are also valid kernels. In this work, we consider the following two pooling schemes:
$\kappa_{+}(\mathcal{G}_{1},\mathcal{G}_{2})$ $=\sum_{\gamma=1}^{N_{t}}\alpha_{\gamma}\,\kappa_{\gamma}(\mathcal{G}_{1},\mathcal{G}_{2}),$ (19)
$\kappa_{\times}(\mathcal{G}_{1},\mathcal{G}_{2})$ $=\prod_{\gamma=1}^{N_{t}}\kappa_{\gamma}(\mathcal{G}_{1},\mathcal{G}_{2}),$ (20)
where $\alpha_{\gamma}\geq 0$ and $\sum_{\gamma=1}^{N_{t}}\alpha_{\gamma}=1$. From a feature-map perspective, if $\kappa_{\gamma}(\mathcal{G}_{1},\mathcal{G}_{2})=\langle\bm{\chi}_{\gamma}(\mathcal{G}_{1}),\bm{\chi}_{\gamma}(\mathcal{G}_{2})\rangle$, then Eq. (19) corresponds to the concatenated map
$\bm{\chi}_{+}(\mathcal{G})=\big\|_{\gamma=1}^{N_{t}}\left(\sqrt{\alpha_{\gamma}}\,\bm{\chi}_{\gamma}(\mathcal{G})\right),$ (21)
where $||_{\gamma=1}^{N_{t}}$ denotes concatenation, while Eq. (20) corresponds to the tensor-product map
$\bm{\chi}_{\times}(\mathcal{G})=\bigotimes_{\gamma=1}^{N_{t}}\bm{\chi}_{\gamma}(\mathcal{G}),$ (22)
where $\otimes$ denotes the tensor product. These constructions suggest that pooling can improve expressiveness, but this is not guaranteed and depends on the additional kernels. We formalize this intuition in Proposition 4 (see App. F for a proof).
@@ III.4 Kernel pooling
Let $S=\{\kappa_{1},\dots,\kappa_{N}\}$ be a set of graph kernels, and let $\kappa_{A}$ and $\kappa_{B}$ be kernels obtained by pooling two subsets $S_{A}$ and $S_{B}$ of $S$ using (19) or (20) as pooling rules. If $S_{A}\subset S_{B}$, then $\kappa_{B}$ is at least as expressive as $\kappa_{A}$.
@@
Classical Random Walk 85.60 $\pm$ 4.23 58.12 $\pm$ 4.76
Graphlet sampling 84.76 $\pm$ 5.41 60.69 $\pm$ 5.11
$k$-core 83.58 $\pm$ 6.00 58.98 $\pm$ 4.66
Shortest path 82.25 $\pm$ 5.69 54.06 $\pm$ 5.12
WL optimal assignment 86.88 $\pm$ 5.01 58.07 $\pm$ 5.00
[…]
$\text{Shortest path}^{\mathrm{(attr)}}$ 81.70 $\pm$ 5.40 59.71 $\pm$ 4.63
$\text{WL optimal assignment}^{\mathrm{(attr)}}$ 86.63 $\pm$ 5.25 63.40 $\pm$ 4.23
Quantum features $\mathrm{QEK}^{\mathrm{(loc)}}$ 86.19 $\pm$ 5.03 64.11 $\pm$ 4.75
$\mathrm{QEK}^{\mathrm{(glob)}}$ 85.80 $\pm$ 5.81 60.62 $\pm$ 5.16
$\mathrm{GDQC}^{\mathrm{(loc)}}$ 89.29 $\pm$ 4.89 60.87 $\pm$ 5.49
$\mathrm{GDQC}^{\mathrm{(glob)}}$ 88.53 $\pm$ 4.33 59.61 $\pm$ 5.57
$\mathrm{QEK}^{\mathrm{(loc)}}_{\bm{+}}$ 89.26 $\pm$ 5.40 69.55 $\pm$ 4.98
$\mathrm{QEK}^{\mathrm{(loc)}}_{\bm{\times}}$ 88.78 $\pm$ 4.89 69.28 $\pm$ 5.28
$\mathrm{QEK}^{\mathrm{(glob)}}_{\bm{+}}$ 89.45 $\pm$ 4.02 62.22 $\pm$ 5.54
$\mathrm{QEK}^{\mathrm{(glob)}}_{\bm{\times}}$ 88.96 $\pm$ 4.98 63.35 $\pm$ 4.78
$\mathrm{GDQC}^{\mathrm{(loc)}}_{\bm{+}}$ 89.35 $\pm$ 4.97 65.86 $\pm$ 5.26
$\mathrm{GDQC}^{\mathrm{(loc)}}_{\bm{\times}}$ 90.83 $\pm$ 5.56 66.51 $\pm$ 4.82
$\mathrm{GDQC}^{\mathrm{(glob)}}_{\bm{+}}$ 88.84 $\pm$ 5.42 59.82 $\pm$ 6.02
$\mathrm{GDQC}^{\mathrm{(glob)}}_{\bm{\times}}$ 89.10 $\pm$ 5.55 61.18 $\pm$ 5.66
@@ IV Results
In this section we present the extensive numerical results where we compare the various classical and quantum-feature kernel implementations introduced in Sec. III with respect to their prediction performance using a support vector machine. Table 1 summarizes the results, providing a full comparison across all models and their respective best performances.
We begin with a brief summary of the numerical-experiment setup. As classical baselines, we consider five widely used graph kernels (cf. Subsec. III.1): the random-walk kernel, the graphlet (sampling) kernel [38], the $k$-core kernel [33, 45], the shortest-path kernel, and the Weisfeiler-Lehman optimal assignment kernel [27]. The last three baselines can be used either with or without node features; when node features are available, we denote the corresponding variant by the superscript $\mathrm{(attr)}$. Classical kernel baselines are computed using the GraKeL library [49]. The two quantum-feature kernels central to this study are the quantum evolution kernel (Subsec. III.2), which is based on the global excitation-probability observable $P_{k}^{\mathcal{G}}(t)$ (6), and the generalized-distance quantum-correlation kernel (Subsec. III.3), which is based on a local observable: the correlation matrix $C^{\mathcal{G}}_{ij}(t)$ (6).
@@
Throughout this work, we train an SVM classifier using precomputed kernel matrices (see the beginning of Sec. III for details on kernel SVMs). We evaluate performance using the weighted F1 score, which accounts for class imbalance (notably in MUTAG and PTC_FM). Concretely, the F1 score is computed per class and averaged with weights proportional to the class support. We employ repeated stratified 5-fold cross-validation with 10 repetitions and report the mean and standard deviation of the weighted F1 score across all folds and repetitions. For each kernel, we tune the SVM regularization parameter $C$ over 100 logarithmically spaced values in $10^{-3},10^{2}]$ and retain the best-performing model. The best results are summarized in Tab. 1. Overall, incorporating node features improves performance. The best classical baseline is the WL optimal assignment kernel: with node features on PTC_FM, and on unlabeled graphs for MUTAG.
@@
Figure 3: Weighted F1 scores associated with SVMs trained on the quantum-feature kernels, $\mathrm{QEK}$ (15) and $\mathrm{GDQC}$ (17), for the two datasets, (a) MUTAG and (b) PTC_FM. The superscripts $\mathrm{(loc)}$ and $\mathrm{(glob)}$ refer to whether the governing Hamiltonian has node features embedded $\hat{\mathcal{H}}_{\mathrm{loc}}^{\mathcal{G}}$ (10)] or not $\hat{\mathcal{H}}_{\mathrm{glob}}^{\mathcal{G}}$ (9)]. The training and evaluation procedure is separately conducted at each time steps across the time evolution. For comparison, the corresponding randomly stratified baseline, where each class is predicted proportionally to its frequency in the dataset are 0.56 for MUTAG and 0.53 for PTC_FM, significantly below the presented F1 scores. The $\mathrm{GDQC}$ results are derived for a number of bins $N_{\mathrm{bins}}^{C}=10$. The light blue and red horizontal dashed lines correspond to the classical early-time baselines for $\mathrm{QEK}$ [see Eq. (16)] and $\mathrm{GDQC}$ [see Eq. (18)], respectively.
Next, we focus on the quantum-feature kernels. Observables are sampled at $k=20$ equally spaced time steps $\{t_{1},t_{2},\dots,t_{k}=t_{\mathrm{max}}\}$. We consider two time-evolution durations: a shorter evolution with $t_{\mathrm{max}}=1\,\mu\mathrm{s}$ and a longer evolution with $t_{\mathrm{max}}=2\,\mu\mathrm{s}$. We begin with the non-pooled results. At each time step, a quantum-feature kernel matrix, $\mathrm{QEK}^{(\mathrm{loc}/\mathrm{glob})}$ or $\mathrm{GDQC}^{(\mathrm{loc}/\mathrm{glob})}$, is precomputed and used to train a SVM. The resulting weighted F1 scores for $t_{\mathrm{max}}=1\,\mu\mathrm{s}$ on both datasets (MUTAG and PTC_FM) are shown in Fig. 3. Overall, the quantum-feature kernels substantially outperform the random baseline and, for their best-performing variants, exceed the strongest classical baselines on both datasets (Tab. 1).
The horizontal dashed lines in Fig. 1 show the classical baselines obtained by evaluating the short-time perturbative expansions of the quantum features at a fixed time (Subsec. II.4): red for the $\mathrm{GDQC}$ baseline (cf. Eq. (18)) and light blue for the $\mathrm{QEK}$ baseline (cf. Eq. (16)). At short times, $\mathrm{GDQC}$ behaves consistently with its early-time baseline. Only once non-negligible correlation elements have built up does the $\mathrm{GDQC}$ performance depart from this baseline. By contrast, particularly on MUTAG, the $\mathrm{QEK}$ kernel shows slightly weaker agreement with its early-time baseline. This discrepancy may stem from multiple factors, including finite-shot noise in the estimation of the observables and the coarse nature of the early-time approximation. In particular, already at the smallest sampled time step $t=40\,\mathrm{ns}$, the largest graphs in the dataset exhibit a non-negligible $P_{2}(t)$ contribution.
More generally, $\mathrm{QEK}$ and $\mathrm{GDQC}$ are not consistently comparable across all settings. In particular, $\mathrm{GDQC}$ attains the best score on MUTAG, whereas $\mathrm{QEK}$ achieves the best score on PTC_FM (Tab. 1). Regarding the use of the node-feature embedded Hamiltonian $\hat{\mathcal{H}}_{\mathrm{loc}}^{\mathcal{G}}$ versus the global-detuning Hamiltonian $\hat{\mathcal{H}}_{\mathrm{glob}}^{\mathcal{G}}$, we observe an overall trend in favor of local detuning across both datasets, with a particularly pronounced effect on PTC_FM (Tab. 1). This pattern is consistent with the classical setting, where attributed variants often outperform their non-attributed counterparts. As the sampling time varies, performance changes only moderately, within $\sim 5$$10\%$ in weighted F1 score, see Fig. 3. The best-performing time step (i.e., the maximum over the sampled times) is reported in Tab. 1; note that it can originate from the time interval $1\,\mathrm{\mu s}<t\leq 2\,\mathrm{\mu s}$, which is not shown in Fig. 3.
Next, we consider pooling schemes to further enhance the quantum-feature kernels. The two pooling schemes are described in Subsec. III.4. Since the choice of which of the $k=20$ time steps to pool (and how many to pool) is a priori arbitrary, we fix the tuple size and evaluate all candidate time-step tuples considered in our sweep. The goal is to quantify the performance gains achievable through pooling and to assess the sensitivity of the models to the pooling choice. For each tuple size, we select the pooling configuration that maximizes cross-validated performance and report the corresponding test performance. Focusing on tuple sizes $3$ and $5$, Tab. 2 summarizes the selected pooled models for $t_{\mathrm{max}}=1\,\mu\mathrm{s}$ and $t_{\mathrm{max}}=2\,\mu\mathrm{s}$, together with the chosen time steps.
Pooling can significantly improve performance: adding the best pooled models from each kernel class to Tab. 1 yields an F1-score increase of $\sim 1\%$ on MUTAG and $\sim 5\%$ on PTC_FM. Overall, the best-performing quantum-feature models outperform the classical baselines. Interestingly, the ranking of kernels is markedly dataset dependent: configurations that perform strongly on MUTAG do not necessarily remain top performers on PTC_FM, and vice versa. Consistent with this observation, $\mathrm{GDQC}$ achieves the best performance on MUTAG, whereas $\mathrm{QEK}$ achieves the best performance on PTC_FM once pooling is applied (Tab. 1). A similar dataset dependence is also visible for classical baselines such as the random-walk and shortest-path kernels (Tabs. 1-2). For the quantum-feature kernels in particular, the choice of representation (local vs. global observables) and the pooling strategy can further shift the ordering (Tab. 2). We therefore treat these design choices as standard model-selection hyper-parameters to be selected via validation.